OpenStudy (anonymous):

i dont get this question ( Contrast roles of the coefficients and exponents when multiplying and dividing terms.)

OpenStudy (kropot72):

An example might help you. What actions do you take with the coefficients of the terms and with the exponents in this product: \[\large 3y ^{2} \times 7y ^{5}=\ ?\]

OpenStudy (anonymous):

you combine like terms

OpenStudy (kropot72):

So to find the product, what operation do you do to the coefficients 3 and 7, that is do you add, divide or multiply.

OpenStudy (anonymous):

add

OpenStudy (anonymous):

right?

OpenStudy (kropot72):

Not really. What is the result of: \[\large 3\times7=\ ?\]

OpenStudy (anonymous):

21

OpenStudy (kropot72):

Correct. So that is what we do to the coefficients 3 and 7 of \[\large 3y ^{2} \times 7y ^{5}\]

OpenStudy (anonymous):

oh i get it its the exponent i add right

OpenStudy (anonymous):

21^7 ?

OpenStudy (kropot72):

What happened to y?

OpenStudy (anonymous):

idk the comment left when i was trying to answer it

OpenStudy (anonymous):

i thought you deleted it or something

OpenStudy (anonymous):

well it would have been 4 and then id have to minus the exponents ?

OpenStudy (kropot72):

Sorry. I want to finish multiplication first. So in your result 21^7 what happened to y.

OpenStudy (anonymous):

oh yeah 21y^7

OpenStudy (kropot72):

Correct! Now lets look at division of terms. What do you do to coefficients and to exponents in: \[\large \frac{16x ^{6}}{4x ^{2}}\]

OpenStudy (anonymous):

so You add the exponents when its multiplying and minus when its dividing ? and 4x^4

OpenStudy (kropot72):

You are correct. Can you answer the question that you posted now?

OpenStudy (anonymous):

yes thank you :) can you help me with another one?

OpenStudy (kropot72):

Sorry, I have to log out now. Please post the other question as a new one.